Multiply.
(4xy3) ( 2x2y )( 5xy2)

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

Multiply.
(4xy^3) ( 2x^2y )( 5xy^2)
Multiply the coefficients to get: 40
Multiply the x-factors to get: x^4
Multiply the y-factors to get: y^6
-----------------
Answer: 40x^4y^6
=====================
Cheers,
Stan H.

Answer by checkley71(8403)

(4XY^3)(2X^2Y)(5XY^2)=(4*2*5)(X*X^2*X)(Y^3*Y*Y^2)=40X^4Y^6 ANSWER.